# Affordance-Aware Object Insertion via Mask-Aware Dual Diffusion

### v1

@@ 3 Dataset
We constructed the SAM-FB benchmark dataset for the affordance-aware object insertion task.
Design Decisions. We designed the dataset to have: 1) Sufficient training samples. 2) A wide variety of foreground objects. 3) Well-aligned input-output pairs such as $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$ tuples, where $\boldsymbol{f}$, $\boldsymbol{b}$, $\boldsymbol{p}$, and $\boldsymbol{x}$ represent the foreground image, background image, position prompt, and ground truth image, respectively. Unfortunately, there is no existing dataset that meets the aforementioned requirements. As shown in Table 1, existing datasets such as DreamEditBench [26] and MureCom [32] contain only a very limited number of samples and object categories.
Dataset Sample No. Category No.
DreamEditBench 440 22
MureCom 640 32
SAM-FB (Ours) 3,160,403 3,439
Table 1: Dataset comparison. Our dataset contains significantly more training samples and object categories.
Our dataset is built upon the SA-1B [23] dataset to ensure a diverse category of foreground objects. SA-1B contains 11 million high-quality images collected from all kinds of reality scenes and 1 billion well-annotated class-agnostic masks. We masked out the foreground object and inpainted the mask to obtain the background. The ground truth is the original image. To automate the process and improve the dataset quality, a scalable automatic annotation pipeline, and a data quality control are proposed. In the end, we obtain over 3 million well-annotated training samples organized in $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$ format, where $\boldsymbol{p}$ consists of diverse position prompts generated from foreground masks.
Figure 1: Pipeline of constructing the SAM-FB dataset. The background is inpainted and high-quality foreground objects are preserved through a data quality control stage.
Figure 2: The framework of MADD. Foreground objects are encoded using a DINOv2 encoder, serving as the guidance signal through the cross-attention mechanism. The position prompt encoder unifies different types of position prompts, which are then concatenated with the latent mask $\mathbf{m_{t}}$. The background is encoded using a VAE encoder and then concatenated with the latent image $\mathbf{z_{t}}$. We use a dual branch structure to denoise RGB image $\mathbf{z}$ and object mask $\mathbf{m}$ simultaneously.
Scalable Automatic Annotation Pipeline. As shown in Figure 1, the pipeline generates foreground-background pairs from images in the SA-1B dataset. For each input image, we first run SAM to obtain object masks as candidates. We then perform the Non-Maximum Suppression (NMS) with a high threshold (0.6) to deduplicate these masks. With these masks, we crop the objects to form the foreground $\boldsymbol{f}$, and the remaining content is inpainted with LAMA [43] to create the background $\boldsymbol{b}$. To ensure the inpainting quality of the background, we expand the boundaries of the object mask to avoid residual foreground content in the background following GroundingSAM [30, 38]. The original image serves as the ground truth $\boldsymbol{x}$. The binary object mask is designated as $\boldsymbol{m}$ and we can derive multiple different positions prompts $\boldsymbol{p}$ such as point, bounding boxes using it, so that compose the input-output pairs $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$, This pipeline is scalable and no manual annotations are required in this process. With this pipeline, it’s feasible and easy for users to further scale up the SAM-FB dataset or customize their datasets with new source images.
Data Quality Control. We only keep high-quality foregrounds, since we may obtain masks that include unintended elements due to challenges in controlling the granularity and semantics of the segmentation while using SAM. They typically suffer from three issues: 1) Incompleteness. Some objects are obscured by others, resulting in only a small part of the original objects being visible. 2) Background Masks. It is also possible to get some background masks such as a piece of sky or a coast, which cannot be treated as a foreground object. 3) Undersized Objects. Some tiny objects feature blurry details, thereby reducing data quality. To achieve quality control, we employ a combination of rule-based and learning-based methods to filter the foreground. We first filter masks by setting rules on their basic attributes. Then, we further employ a deep neural network to filter out low-quality objects. We annotate 2,000 foreground images with binary labels (good and bad) and fine-tune a pre-trained ResNet-50 [15] on those manually annotated images. The model is then able to assess the quality of the foreground. After the strict quality control, only 0.25% of the masks are left. Samples of low-quality objects and details of the filters are in the supplementary.
To justify the diversity of our proposed dataset, we use RAM [51] to recognize the categories of foregrounds we keep. We have identified 3439 different categories, which is a vast improvement over previous datasets. The word cloud of categories is shown in our supplementary.
4 Method
4.1 Task Formulation
We formally define the affordance-aware object insertion problem as follows: given a foreground object $\boldsymbol{f}$, a background scene $\boldsymbol{b}$, and a position prompt $\boldsymbol{p}$, the model $\mathcal{G}$ is required to predict a synthesized image $\hat{\boldsymbol{x}}$ that places the object into the scene following the correct affordances and aligns with the intentions of the position prompt:
$\hat{\boldsymbol{x}}=\mathcal{G}(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p}).$ (1)
The model must support flexible prompts $\boldsymbol{p}$, which include points, bounding boxes, masks, and even null prompts.
4.2 Mask-Aware Dual Diffusion
Inspired by the recent success of diffusion models [39], we adopt SD as the backbone of our model to achieve high-quality image synthesis. SD with condition guidance represents a Markov process where Gaussian noise $\boldsymbol{\epsilon}_{t}\sim\mathcal{N}(0,1)$ is added to the ground truth $\mathbf{z}_{0}$ to produce $\mathbf{z}_{t}=\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t}$. Here, $\alpha_{t}$ and $\sigma_{t}$ are scalars sampled by different samplers such as DDIM [41] and DDPM [17]. The objective of the SD is to reconstruct the reverse process $p_{\theta}(\mathbf{z}_{0}|\mathbf{c})$:
$p_{\theta}(\mathbf{z}_{0:T}|\mathbf{c})=p(\mathbf{z}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{c}),$ (2)
where $\mathbf{c}$ is the condition to guide the denoising process. In our task, $\mathbf{z}_{0}=\boldsymbol{x}$ and $\mathbf{c}$ consists of a foreground image $\boldsymbol{f}$, a background image $\boldsymbol{b}$, and a position prompt $\boldsymbol{p}$.
For affordance-aware object insertion, one key goal is to determine an appropriate location and size for the object to be inserted. The final inserted region, e.g., the insertion mask $\boldsymbol{s}$ contains the accurate location, size, and shape information, yet hasn’t been explicitly modeled in Eq. 2. To facilitate the learning of the affordance concept, we explicitly model the insertion mask $\boldsymbol{s}$ in the diffusion process. Specifically, we propose a Mask-Aware Dual Diffusion model, which adopts a dual-stream architecture to simultaneously denoise the appearance and the insertion mask. Instead of just predicting the composed image, MADD also predicts the precise mask $\hat{\mathbf{m}}$ of the desired inserted object explicitly at each time step. This means that the reverse process of $\mathbf{m}_{t}$ and $\mathbf{z}_{t}$ also depends on each other, so we have $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{m}_{t},\boldsymbol{f},\boldsymbol{b},\boldsymbol{p})$ for the RGB stream and $p_{\theta}(\mathbf{m}_{t-1}|\mathbf{m}_{t},\mathbf{z}_{t},\boldsymbol{f},\boldsymbol{b},\boldsymbol{p})$ for the mask stream, where $\mathbf{m}_{t}=\alpha_{t}\mathbf{m}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t}^{m}$, $\boldsymbol{\epsilon}_{t}^{m}\sim\mathcal{N}(0,1)$, and $\mathbf{m}_{0}=\boldsymbol{s}$. Note that although $\boldsymbol{s}$ is included in our SAM-FB dataset, it can be directly derived from the difference between the ground truth $\boldsymbol{x}$ and the background $\boldsymbol{b}$.
For the diffusion process of MADD, we have
$p_{\theta_{1}}(\mathbf{z}_{0:T}|\mathbf{c},\mathbf{m}_{0:T})=p(\mathbf{z}_{T})\prod_{t=1}^{T}p_{\theta_{1}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{m}_{t},\mathbf{c})$ (3)
$p_{\theta_{2}}(\mathbf{m}_{0:T}|\mathbf{c},\mathbf{z}_{0:T})=p(\mathbf{m}_{T})\prod_{t=1}^{T}p_{\theta_{2}}(\mathbf{m}_{t-1}|\mathbf{z}_{t},\mathbf{m}_{t},\mathbf{c}),$
where $\theta_{1}$ and $\theta_{2}$ represents parameters for different tasks: $\theta_{i}=\theta_{shared}\cup\theta_{expertise}^{i}$. Every time step, the denoiser $\boldsymbol{\epsilon}_{\theta}$ takes the foreground, background, position prompt, and time step $t$ as input to predict the noise map $\boldsymbol{\epsilon}_{t}$ for the RGB image and $\boldsymbol{\epsilon}_{t}^{m}$ for the mask as follows:
$\boldsymbol{\epsilon}_{t},\boldsymbol{\epsilon}_{t}^{m}=\boldsymbol{\epsilon}_{\theta}(\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t},\alpha_{t}\mathbf{m}_{0}+\sigma_{t}\boldsymbol{\epsilon}^{m}_{t},\mathbf{c}).$ (4)
Figure 2 shows the framework of the MADD. Inspired by Liu et al. [31], we utilize a single UNet with two expert input-output branches to denoise them simultaneously. The two tasks share the entire UNet except for the conv_in, first Down Block, last Up Block, and the conv_out. These independent blocks serve as the expertise input-output branch. Skip connection is also performed between the corresponding expertise input and output branches. This dual diffusion architecture reuses most of the parameters, encouraging the model to incorporate knowledge from two relevant tasks.
Modeling $p_{\theta_{1}}$ and $p_{\theta_{2}}$ is equivalent to estimating the noise terms $\boldsymbol{\epsilon}_{t}$ and $\boldsymbol{\epsilon}_{t}^{m}$ at each time step. This can be achieved by minimizing the loss function for RGB images and insertion masks:
$\mathcal{L}_{z}(\theta_{1})=\mathbb{E}\left[w_{t}\|\boldsymbol{\epsilon}_{t}-\boldsymbol{\epsilon}_{\theta}(\alpha_{t}\mathbf{z}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t};t,\mathbf{m_{t}},\mathbf{c})\|^{2}\right]$ (5)
$\mathcal{L}_{m}(\theta_{2})=\mathbb{E}\left[w_{t}\|\boldsymbol{\epsilon}_{t}^{m}-\boldsymbol{\epsilon}_{\theta}(\alpha_{t}\mathbf{m}_{0}+\sigma_{t}\boldsymbol{\epsilon}_{t}^{m};t,\mathbf{z_{t}},\mathbf{c})\|^{2}\right]$ (6)
$w_{t}$ is a function weighing the contributions of denoising tasks to the training objective, commonly set to 1. During the training procedure, the time step $t$ is sampled from a uniform distribution $\mathcal{U}[0,1]$. The final loss function for MADD is obtained by combining the above losses:
$\mathcal{L}(\theta)=\mathcal{L}_{z}(\theta_{1})+\mathcal{L}_{m}(\theta_{2}).$ (7)
4.3 Affordance Condition
Foreground Encoder. In MADD, foreground images serve as a guidance condition and are injected into the mainstream using cross-attention. The foreground embedding can be any tokenized features obtained by a ViT-like visual encoder. In SD, the CLIP [37] encoder is used to provide text semantics for cross-attention. Following Efficient-3DiM [20], we use DINOv2 [35] as the encoder. While CLIP visual features are designed to align with textual features, typically encapsulating high-level semantic features, DINOv2 features preserve detailed object information, which is crucial for our task.
Background Encoder. Following practices in SD-based image editing methods [42, 5], we use a frozen pre-trained VAE Encoder to encode the background image to a 4-channel latent map. As the background image is pixel-aligned with the output, we concatenate the background latent map with the $\mathbf{z_{t}}$ and send them to the UNet. The pre-trained VAE decoder is then used to transform the latent output back to RGB space.
Position Prompt Encoder. To accommodate various position prompts, we have designed a unified representation that processes different types of prompts. Specifically, we convert both sparse and dense position prompts into a unified dense representation. Each type of prompt is represented as a 1-channel position map. Note that these position prompts serve as rough guides for the placement of foreground objects, therefore we consider building a probability map for the 1-channel position map. For point prompts, we transform them into a Gaussian heatmap where the point is positioned at the center of the Gaussian. When dealing with bounding box prompts, we fill the area inside the box with ones and the outside with zeros. For mask prompts, we set the values inside the mask to 1 and all other values to 0. For the null prompt, we use an all-one image, implying that all positions are possible. Subsequently, we resize the position map to match the spatial dimensions of the latent map $\mathbf{m}_{t}$, and concatenate it with $\mathbf{m}_{t}$ along the channel dimension before feeding it into the diffusion U-Net.
4.4 Implementation Details
To leverage prior knowledge of real-world images, we initialize our model with a pre-trained Stable Diffusion Inpainting model for the neural layers with the same configuration. Fine-tuning a diffusion model is resource-consuming, but we can make it more efficient using the feature that the model is image-size unaware. Therefore, we first train the model on $128\times 128$ resolution to accelerate training, with a batch size of 1024 on 2 A-100 GPUs for 35K steps. We then fine-tune it on $256\times 256$ resolution with a batch size of 256 for 15K steps. For each foreground object, all different position prompts $\mathbf{p}$ can be generated with the mask $\mathbf{s}$. During training, we randomly selected one of the position prompts as input with equal probability. We use Classifier-Free Guidance [16] to improve the insertion quality. During training, we drop all the conditions with a probability of 0.1 and only drop $\mathbf{p}$ with a probability of 0.1. We also perform strong data augmentation to both foreground images and position prompts to prevent the model from learning a mere copy-and-paste process. Please refer to the supplementary for details.
5 Experiments
5.1 Results on the SAM-FB Test Set
Evaluation Metrics. Following the previous work [33, 19, 22, 6], we employ the FID score to measure the quality of images obtained by generative models and the CLIP score to evaluate the semantic similarity between the edited region and the reference foreground.
(a) Location Adjustment
(b) View Adjustment
(c) Size Adjustment
(d) Automatic Localization
Figure 4: We test ambiguous prompts (points and blank) on the in-the-wild images. When providing the prompt of point, 4(a), 4(b), and 4(c) show that our model can adjust properties of foreground objects to achieve the affordance insertion.4(d) illustrates that the model could find the suitable position to insert the object.
Quantitative Comparison. We compare our method with Stable Diffusion [39], PBE [47], GLI-GEN [29], and Human Affordance [24] on the SAM-FB test set to show the effectiveness of our method. The results in Table 2 show that our method attains the best FID score and the highest CLIP score, illustrating the superiority of our model. We further present our results with different prompts in Table 3, we see that the mask prompt achieves the best results as it provides more accurate position information.
Qualitative Results. We left 3786 images as test split. Figure 3 presents the visualization results on the SAM-FB test set. In each group, the leftmost image depicts the background marked with a position prompt. Our MADD predicts the RGB image and mask of the inserted object, which are shown in the last two images of each group. These results demonstrate that MADD not only inserts objects with high quality but also accurately predicts object masks. Figure LABEL:fig:teaser presents additional results from the SAM-FB test set, illustrating that our model is capable of performing affordance-aware object insertion.
[…]
mask bbox mask bbox mask bbox
 [39] 15.41 15.47 0.7079 0.8058 860 883
 [47] 33.68 24.59 0.7664 2373 1615
 [29] 14.21 0.7944 830
 [24] 14.49 14.42 0.8014 0.8637 857 845
Ours 13.53 13.60 0.8727 0.8658 760 775
Prompt Mask Bbox Point Null Avg.
FID 13.53 13.60 13.66 13.96 13.69
MSE 760 775 772 860 792
CLIP Score 0.8727 0.8658 0.8567 0.8034 0.8415
Table 3: Comparison of position prompts on the SAM-FB test set.
Ablation Study. Table 4 presents the results of our ablation study. We replace the image encoder in the Human Affordance model with DINOv2 as the baseline. Next, we add Classifier-Free Guidance to enhance the affordance condition signal. Then we diffuse RGB image and object mask simultaneously, but sharing the entire UNet. Finally, we use two expertise branches for mask and RGB streams. The results show that all of them improved performance.
Method FID ($\downarrow$) CLIP×100 ($\uparrow$)
Baseline 25.89 89.12
+ Classifier-Free 21.93 91.13
+ Dual diffusion 21.75 91.57
+ Expertise branch 21.55 91.68
Table 4: Ablation study on the SAM-FB test set with $128\times 128$ resolution using mask prompts.
5.2 Results on In-the-wild Images
Ambiguous Prompts. The training process with position augmentations ensures that the model can understand the affordance relationship between the scene and objects. By using MADD, the model explicitly refines the object’s position, making it capable of handling ambiguous position information. Moreover, the model can adjust the object’s position, size, and view to ensure coherence with the background scene. In Figure 4(a), the model adjusts the person’s position around the provided point, placing them on the ground and in front of the bike instead of in the air. Figure 4(b) demonstrates that the model can change an object’s view. The car’s orientation is adjusted to align with the lane. Figure 4(c) shows that the model adjusts the size of the coffee beans to match the background scene, even though only a point prompt and a reference image of the statue are provided. Furthermore, the model is capable of placing objects even in the absence of position prompts, as shown in Figure 4(d). In this case, the model automatically places the cake onto the plate without any additional clues. When no prompts are provided, the model can still determine a feasible insertion. Additionally, when ambiguous prompts are given, MADD can generate diverse and feasible insertions, as shown in Fig. 5. For a prompt of a point, the model finds reasonable insertions around the point, and for a blank prompt, the model searches for suitable insertions globally.
(a) Rank distribution for different methods. Our method has the highest proportion of rank 1 and the least proportion of rank 5.
(b) Rank-1 distribution for each criterion. Each pie chart represents the proportion of times each model achieved Rank-1 for a specific evaluation criterion. Our method dominates every metric.
Figure 6: Human evaluation results on in-the-wild Images. We compared 10 groups of images according to different criteria. Our MADD model outperforms SDXL [39], GLI-GEN [29], ObjectStitch [42] and PBE [47] on overall ranking and each criteria.
Figure 7: MADD can work on images of higher resolution, generating sharper edges, clearer reflections, and improved texture details.
Figure 8: More in-the-wild examples with null prompts. The model can generate an affordance-feasible solution to insert the foreground objects according to the background scene.
Human Evaluation. To test the generalization ability of our MADD model, we performed affordance insertion on in-the-wild images and compared the results with Stable Diffusion XL [36], GLI-GEN [29], PBE [47], and ObjectStitch [42]. Instead of merely relying on metrics like FID and CLIP Score, we also conducted a user study to achieve a more comprehensive comparison.
We asked 10 users to rank 10 groups of composited images generated by different models according to the following criteria: 1) Foreground and Background Integration, 2) Foreground Clarity and Detail, 3) Foreground Appearance Consistency with Reference, 4) Lighting and Shade on Foreground, 5) Color Consistency.
Figure 6(a) shows the distribution of ranks for different models, where rank 1 and rank 5 represent the best and worst quality, respectively. Our model achieves 50% of Rank-1 placements and 1.60% of Rank-5 placements, outperforming other methods. Figure 6(b) provides details for each criterion. Our model consistently achieved the highest proportion of Rank-1 placements across all evaluation criteria, as indicated by the largest segments in each pie chart, especially for maintaining the consistency of foreground appearance with the reference image. This dominance in Rank-1 distribution across multiple criteria highlights the model’s superior performance compared to others in each assessed aspect.
Figure 8 shows additional in-the-wild examples with null prompts. The results demonstrate the strong generalization ability of our method. The supplementary material contains more samples to further illustrate that our model can provide suitable and diverse solutions in in-the-wild scenarios.
5.3 Results on Higher-Resolution Images
Even though we trained our model at a resolution of 256x256, it is also capable of handling higher-resolution inputs, such as 512x512. To demonstrate this, we fine-tuned our model on SAM-FB at a resolution of 512x512 for only 200 steps and compared it with the original checkpoint on the test split. Figure 7 shows the results of the original checkpoint and the fine-tuned version at higher resolution using the same input. We highlighted the details of the inserted object. It is demonstrated that the model generates sharper edges, clearer reflections, and improved texture details after fine-tuning at a higher resolution.
6 Conclusion
In this paper, we extend the notion of affordance to a more generalized setting, rather than being limited to human-centered tasks. We proposed the affordance-aware object insertion task which involves placing any object into any scene using various types of position prompts. To support this challenging task, we have constructed a SAM-FB dataset which consists of 3,160,403 high-quality foreground-background pairs across over 3,000 object categories. We present a dual-stream diffusion model, MADD, which is designed to exploit affordance relationships between foregrounds and backgrounds by simultaneously denoising the foreground appearance and the object mask. Leveraging its understanding of affordance, MADD supports a variety of position prompts, varying from points to masks, and can automatically insert objects even in the absence of explicit prompts. Moreover, MADD exhibits strong generalization capabilities with foreground objects. Our method outperforms previous diffusion-based image composition models on both the SAM-FB test set and in-the-wild images, seamlessly integrating various objects into any background while keeping semantic consistency.
Supplementary Material
Appendix A More Implementation Details
A.1 SAM-FB Dataset
Figure 10(a) shows the object masks before and after the data quality control. We see that with our designed data quality control, the foreground object masks have better quality. Figure 10(b) illustrates the word cloud of our SAM-FB dataset, we observe that our SAM-FB dataset consists of diverse object categories. Table 5 demonstrates the specific threshold for each filter operation of our data quality control stage. With these filter operations, only 0.25% of the masks are left, which ensures the quality of the obtained dataset.
A.2 Traning Details
We constructed our dual-stream UNet based on the Stable Diffusion Inpainting v1.5 model, incorporating several modifications. Specifically, we replicated both the first down-sampling block (including conv_in and the first Down Block) and the last up-sampling block (including the last Up Block and conv_out) to accommodate dual-stream inputs and outputs. To bypass the need to start from scratch, we initialized our model using the pre-trained Stable Diffusion Inpainting v1.5 checkpoint for the unchanged blocks. To optimize our computing resources, we employed a gradual scaling-up approach to train our model. Initially, the model was trained at a resolution of $128\times 128$, with a batch size of 1024 and a learning rate of $1.25\times 10^{-4}$. This phase included 5,000 warm-up steps, followed by a constant schedule. We then fine-tuned the model at a higher resolution of $256\times 256$, reducing the batch size to 256 and adjusting the learning rate to $5\times 10^{-5}$. For the diffusion process, the time step is 1000 with a linear noise scheduler. We use 2048 samples of SAM-FB for testing and the rest for training. To ensure a fair comparison with other methods, we re-implemented and re-trained the closest method, Human Affordance, since the authors did not release either the model weights or the full dataset they used. We adopted the diffusion model architecture provided by the authors and trained it using our SAM-FB dataset. Additionally, we replaced the mask input in the original model with our position map to ensure compatibility with the SAM-FB dataset.
A.3 Data Augmentation

### v2

@@ 3 Dataset
Figure 1: Dataset construction pipeline for SAM-FB. The pipeline automatically converts any input image into a tetrad output through four stages, ensuring high-quality foreground and background retention via a rigorous data quality control process.
Instead of manually designing affordance, we use a data-driven approach to extract affordance relationships between the object and background from the large-scale dataset.
3.1 Motivation & Design Decisions
To support affordance-aware object insertion across various objects and scenes, the dataset must provide rich contextual information and extensive coverage and preserve natural affordance cues between objects and their surroundings. Therefore, we designed the dataset with the following key characteristics: 1) Well-aligned input-output pairs. Rather than using techniques like DreamBooth [43] to memorize each object individually, we adopted supervised fine-tuning to directly learn from large-scale object distributions. Consequently, well-aligned input-output pairs, represented as $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$ tuples, are essential. Here, $\boldsymbol{f}$, $\boldsymbol{b}$, $\boldsymbol{p}$, and $\boldsymbol{x}$ denote the foreground image, background image, position prompt, and ground truth image, respectively. 2) Sufficient training samples. Training a model capable of generalizing to diverse scenarios requires a substantial number of training samples. 3) A wide variety of foreground objects. The dataset must encompass a diverse range of foreground objects to ensure that the trained model can generalize across different object categories. However, existing datasets either lack sufficient object diversity or introduce artifacts that disrupt these affordance relationships.
The $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$-formatted data can be derived from multiple datasets relevant to this task, as shown in Table 1. One option is to use datasets designed for Customized Image Composition. Datasets such as DreamEditBench [27] and MureCom [34] are manually captured by placing the same foreground object in different backgrounds under controlled camera settings. However, the cost of constructing such high-quality paired datasets is substantial, resulting in relatively small dataset sizes. Referring to the Image-Guided Object Insertion task, another approach is to leverage large-scale object detection datasets such as COCO [31] and Open-Images [25] to simulate foreground-background pairs. This is achieved by using bounding boxes as position prompts and extracting foreground objects from the original image using masks, treating the remaining portion as the background. However, these datasets are still insufficient to guarantee robust generalization, particularly due to their limited object category diversity. Moreover, masking out the foreground leaves an opaque region in the background, which disrupts affordance relationships. For HumanAffordance [24], they use a vast number of human-centered video to capture the affordance relationship, but the category is human only and the dataset is private. To address these limitations and obtain a sufficient number of training samples while ensuring diversity in foreground objects, we turned to the SA-1B dataset [23], which consists of 11 million high-quality images collected from various real-world scenes and 1 billion well-annotated class-agnostic masks. Based on SA-1B, we constructed the SAM-FB benchmark dataset for the affordance-aware object insertion task.
3.2 Dataset Construction Pipeline
Figure 1 illustrates the pipeline used to construct the aforementioned $(\boldsymbol{f},\boldsymbol{b},\boldsymbol{p},\boldsymbol{x})$ formatted data. The pipeline automatically transforms any input image into a tetrad output. To enhance affordance representation and improve the overall quality of the SAM-FB dataset, the dataset construction process consists of four key stages.
Foreground Mask Generation. For each unannotated input image, we first apply the Segment Anything Model (SAM) to generate object masks as candidates. Since SAM produces masks at different levels of granularity with varying confidence scores, we perform Non-Maximum Suppression (NMS) with a high threshold (0.6) to remove redundant masks and filter out sub-object-level masks. The remaining masks are used to extract candidate foreground objects.